How much does a 0.36 m long copper rod with a 7.7 mm diameter stretch when one end is fixed and the other is pulled with a 1.5 kN force
Answer #1

The Young Modulus for copper wire is

gama = 120*10^(9) Pa = (F/A)/ (delta L / L)

delta L = FL/(A gama) = 1.5*10^3 * 0.36 / (pi * 0.00385^2 * 120*10^9)

= 0.097*10^(-3) m

= 97 * 10^(-6) m = 97 um